# A large data semi-global existence and convergence theorem for vacuum Einstein’s equations

### v3

@@ 1.6. Main Idea of the Proof
$\Big|\sum_{I\leq 2}e^{-T}\!\int_{M(T)}N\,\nabla^{I}\mathfrak{T}\,\nabla^{I}\Sigma\,\mu_{g}\Big|\lesssim e^{-T}\mathcal{O}^{2}.$
The universal damping factor $e^{-T}$ allows one to close the estimate for $\mathcal{F}$:
$\mathcal{F}\lesssim 1+\mathcal{I}_{0}+\mathcal{O}^{2}.$
The appearance of $e^{-T}$ is the essential difference from the $\Lambda=0$ case: when $\Lambda=0$, no such decay is present, forcing $\mathfrak{T}$ to be chosen small to maintain the bootstrap. This recovers the small-data regime treated by [3] (since by the result of Ye [52], smallness of $\mathfrak{T}$ along with the geometric conditions available in this framework would imply the existence of a hyperbolic metric on $M$). In contrast, for $\Lambda>0$, the exponential damping allows one to treat large data, provided the initial time $T_{0}=a\gg 1$.
The importance of $\mathcal{F}$ becomes evident in the highest-order curvature estimates. One encounters Riccati-type nonlinearities of the form
$\sum_{I\leq 3}\sum_{m=0}^{I}\!\!\sum_{J_{1}+J_{2}+J_{3}=m}e^{-T}\!\!\int_{M(T)}\nabla^{J_{1}}N\,\nabla^{J_{2}+1}\Sigma\,\nabla^{J_{3}+I-m}\Sigma\,\nabla^{I+1}\Sigma\,\mu_{g},$
which can be estimated as
$\lesssim e^{-(1+\gamma)T}\mathcal{F}\mathcal{O}^{2}\lesssim e^{-(1+\gamma)T}\mathcal{O}^{4}\lesssim e^{-(1+\gamma)T}\Gamma^{4}.$
Upon integration in time from $T_{0}$ to $T>T_{0}$, the energy contribution is $\lesssim\bigg[e^{-(1+\gamma)T_{0}}-e^{-(1+\gamma)T}\bigg]\Gamma^{4}\lesssim 1$, which closes the $\mathcal{O}$-estimate. In particular, observe that for an arbitrarily large but finite $\mathcal{I}_{0}$, there always exists a large $T_{0}=a$ leading to this estimate since the relative smallness is dictated by $\mathcal{I}^{0}e^{-a/10}<1$ in the main theorem 1.1. The bootstrap assumptions are thereby improved, completing the argument. This explains the idea behind the relative smallness and allowance for large data.
To summarize, the exponential factor $e^{-T}$- a direct consequence of $\Lambda>0$ introduces a universal dissipative mechanism that couples curvature, geometry, and gauge in a self-stabilizing fashion provided that $T_{0}=a\gg 1$. The interaction between $\mathcal{O}$, $\mathcal{F}$, and $\mathcal{N}^{\infty}$ ensures that curvature control enhances geometric regularity, which in turn stabilizes the gauge, feeding back to maintain bounded curvature. Simultaneously, integrating the transport equation
$\frac{\partial g_{ij}}{\partial T}\approx e^{-T}N\Sigma_{ij}-2\!\left(1-\frac{N}{n}\right)g_{ij},$
shows that the metric coefficients evolve smoothly without degenerating Sobolev constants (see propositions 3.13.2, 3.3, and 3.4 in sections 3.4 and 3.5, respectively for detailed proofs). Thus, the combined bootstrap–damping mechanism yields a globally controlled evolution for large initial data in the $\Lambda>0$ regime.
@@
The analytical framework developed herein extends to the case in which the underlying manifold MM admits a positive Yamabe invariant, i.e., $\sigma(M)>0$, which includes in particular the setting with positive cosmological constant $\Lambda>0$. Under an additional spectral condition on the Laplace–Beltrami operator associated to the evolving metric, ensuring the uniqueness of the solution to the elliptic lapse equation, we establish convergence of the metric to one with constant positive scalar curvature.
@@
In our context, the underlying cause of non-uniqueness is distinct and originates from the failure of uniqueness of solutions to the lapse equation, rather than the conformal degeneracy of the space of constant scalar curvature metrics. We emphasize that, even in the presence of convergence, the limiting metric is generally not conformal to the initial metric. For completeness, and to illustrate the applicability of our method in this setting, we state the corresponding convergence result below; the proof proceeds with only minor modifications to the arguments presented in the negative Yamabe case.
@@
(30) $-\Delta_{g}(1-\frac{N}{n})+(|\Sigma|^{2}-\frac{1}{n})(1-\frac{N}{n})=-|\Sigma|^{2},~\Delta_{g}:=g^{ij}\nabla_{i}\nabla_{j}.$
@@
$\Delta_{g}+n\left(|\Sigma|^{2}-\frac{1}{n}\right)$
@@
Theorem 1.2 (Semi-global existence for $\Lambda>0$ and $\sigma(M)>0$).
@@ 1.7. Consequences for Geometrization
An important conceptual motivation underlying the present work concerns the potential role of the Einstein evolution equations in the geometric classification of closed $3$-manifolds. Specifically, we consider the relevance of the long-time behavior of solutions to the Einstein flow in connection with Thurston’s geometrization conjecture. In this discussion, we assume throughout that the initial data satisfy the regularity assumptions stipulated in the main theorem.
A central open problem in mathematical general relativity is whether the Einstein equations on globally hyperbolic, spatially compact cosmological spacetimes can dynamically implement a form of geometrization of the underlying spatial manifold. In particular, we consider the case in which the Cauchy hypersurface $M$ is a closed, oriented, connected $3$-manifold of negative Yamabe type, i.e., with $\sigma(M)\leq 0$.
@@
Let $M$ be a closed, oriented, connected $3$-manifold with $\sigma(M)\leq 0$. A weak geometrization of $M$ is a decomposition
@@
A heuristic explanation for this phenomenon is furnished by comparison with a classical theorem of Cheeger and Gromov [7, 8], which characterizes volume collapse of Riemannian manifolds with bounded curvature. Specifically, if a sequence $\{(M_{j},g_{j})\}$ of compact Riemannian $3$-manifolds exhibits collapse with two-sided curvature bounds, then $M_{j}$ admits an $\mathcal{F}$-structure of positive rank, and hence is a graph manifold for sufficiently large $j$. Thus, any dynamical mechanism aiming to realize geometrization via Einstein evolution must, at minimum, produce volume collapse with bounded curvature in the graph component $G\subset M$ for large times.
@@
Let $B(1)\subset M$ denote a geodesic ball of unit radius (scale $1$ in injectivity radius scale) with respect to the induced metric at CMC time $T\in[T_{0},\infty)$. Then the volume $\operatorname{Vol}(B(1))_{T}$ remains uniformly bounded above by the initial volume $\operatorname{Vol}(B(1))_{T_{0}}$ for all $T\geq T_{0}$.
@@
1.8. Cosmological Implications
Although our analysis is grounded in the geometric analysis of the Einstein field equations with $\Lambda>0$, we briefly comment on an implication of our results for a longstanding open problem in mathematical cosmology. A central hypothesis in physical cosmology is the so-called cosmological principle, which posits that, at sufficiently large spatial scales, the universe is approximately homogeneous and isotropic. This assumption underlies the derivation of the Friedmann–Lemaître–Robertson–Walker (FLRW) class of solutions, in which constant sectional curvature geometries model the spatial slices.
From a rigorous mathematical perspective, a natural question is whether such large-scale homogeneity and isotropy can arise dynamically from generic initial data under the evolution dictated by the Einstein equations. In particular, one may ask whether spatial slices, even if initially far from being homogeneous or isotropic, can converge in an appropriate topology to a constant sectional curvature geometry—e.g., hyperbolic geometry—along the cosmological constant–driven CMC evolution. This idea was conjectured in [33], where it was suggested that the long-time dynamics of the Einstein-$\Lambda$ flow may select the hyperbolic metric as a universal attractor.
Our analysis, based on Theorem 1.1, provides a rigorous counterexample to this conjecture. Indeed, we work with the most general class of closed, connected, oriented three-manifolds of negative Yamabe type, encompassing all topological classes listed in Cases 1–3 of section 1. For such generic topologies, we demonstrate that no form of geometric convergence of the rescaled spatial metric to a hyperbolic (i.e., maximally symmetric) geometry occurs.
Corollary 1.3 (Failure of the Cosmological Principle as a Derived Consequence).
[…]
@@ 2. Preliminaries
$\mathcal{L}_{g}:\Gamma(S^{2}T^{*}M)\to\Gamma(S^{2}T^{*}M),$
@@
Throughout, the spaces of smooth symmetric covariant 2-tensors and vector fields on $M$ are denoted by $S^{0}_{2}(M)$ and $\mathfrak{X}(M)$, respectively. The trace-free and transverse-traceless parts of a symmetric 2-tensor $A$ are denoted by $\widehat{A}$ and $A^{TT}$, respectively. We will also frequently denote the Cauchy slice at constant mean curvature time $T$ by $M(T)$. In addition, we frequently use the following transport identity which is a consequence of the transport theorem [32]
@@ 3. Gauge fixed Einstein’s equations
Initial data for 40 consist of an n dimensional manifold $M$, a Riemannian metric $g$, a covariant 2-tensor $k$, all assumed to be smooth and to satisfy
@@
where $k^{TT}$ denotes the transverse-traceless (TT) part of $k$, satisfying $tr_{g}k^{TT}$ and $\mbox{div }_{g}k^{TT}=0$, and $Z$ is a smooth vector field on $M$.
@@
We show that the non-TT part involving the vector field Z must vanish. To that end, consider the $L^{2}$-inner product of the momentum constraint with the vector field Z, and apply Stokes’ theorem:
@@
(55) $R(g)-|k^{\mathrm{TT}}|^{2}=2\Lambda-\tau^{2}\left(1-\frac{1}{n}\right).$
@@
(69) $\mathfrak{T}_{ij}:=\text{Ric}_{ij}-\frac{1}{n}R(g)g_{ij}.$
@@ 3.1. Transported Spatial Coordinates and the Re-Scaled Einstein System
From (72), and the fact that $\widehat{\mathsfbf{n}}=\frac{1}{N}(\partial_{T^{\prime}}-X^{i}\partial_{i})$ in general coordinates, we deduce:
(73) $\mathscr{L}_{\widehat{\mathsfbf{n}}}(x^{i})=\widehat{\mathsfbf{n}}(x^{i})=\frac{1}{N}(\partial_{T^{\prime}}-X^{j}\partial_{j})(x^{i})=-\frac{1}{N}X^{i}\Rightarrow X^{i}=0.$
@@
(76) $-\Delta_{g}\left(\frac{N}{n}-1\right)+\left(|\Sigma|^{2}+\frac{1}{n}\right)\left(\frac{N}{n}-1\right)=-|\Sigma|^{2}.$
@@ 3.2. Integration and Norms
Let $(M,g)$ be a smooth, closed, oriented Riemannian $3$-manifold. Fix a smooth partition of unity $\{\xi_{U}\}_{U\in\mathcal{U}}$ subordinate to a finite atlas $\{(U,\varphi_{U})\}_{U\in\mathcal{U}}$ of coordinate charts $\varphi_{U}:U\to\mathbb{R}^{3}$. For a measurable function $f:M\to\mathbb{R}$, the integral of $f$ with respect to the Riemannian volume form $\mu_{g}$ is defined by
@@
We also introduce exponentially weighted norms in time to account for the expanding geometry in cosmological spacetimes. Let $T$ denote the CMC time parameter, and let $\gamma>0$ be a scaling parameter (which will ultimately be chosen to be $\gamma=1$). For $I\leq 2$, define the weighted Sobolev norms
@@
where $N$ denotes the spacetime lapse function.

### v4

@@ 1.7. Main idea of the proof
(36) $\partial_{T}\mathfrak{T}$ $=e^{-T}\,\mathcal{Q}_{2}\bigl(g,N,\Sigma,\mathfrak{T},\nabla\Sigma,\nabla^{2}\Sigma,\nabla N,\nabla^{2}N\bigr)+\Bigl(\frac{N}{n}-1\Bigr)g,$
where $\mathcal{Q}_{1}$ and $\mathcal{Q}_{2}$ denote universal linear combinations of tensorial contractions of the indicated quantities and their covariant derivatives. In particular, the two terms that are most problematic in the vacuum case,
$\mathfrak{T}\qquad\text{and}\qquad\Sigma*\Sigma,$
now enter the evolution only through $e^{-T}\mathfrak{T}$ and $e^{-T}\Sigma*\Sigma$. This is the basic structural fact on which the whole argument rests.
We emphasize that this mechanism is not a null structure in the classical sense of Klainerman, since the gain does not arise from an algebraic cancellation in the quadratic form relative to the characteristic cone of the principal hyperbolic operator. Rather, the positive cosmological constant produces a weak-null-type integrable damping mechanism: the nonlinear interactions that would otherwise be borderline are multiplied by a universal coefficient $e^{-T}\in L^{1}([T_{0},\infty))$. This integrability is exactly what makes large–data forward control possible once the initial CMC slice is placed sufficiently far in the expanding regime.
To exploit this structure, we introduce three scale–invariant quantities. The first is the high–order geometric energy
$\mathcal{O}(T):=\sum_{I\leq 3}\Bigl(\|\nabla^{I}\Sigma\|_{L^{2}(M_{T})}+\|\nabla^{I-1}\mathfrak{T}\|_{L^{2}(M_{T})}\Bigr),$
which controls the differentiated curvature and deformation variables. The second is the lower–order renormalized quantity
$\mathcal{F}(T):=\sum_{I\leq 2}\|e^{T}\nabla^{I}\Sigma\|_{L^{2}(M_{T})},$
equivalently
$E^{\mathrm{lower}}(T):=e^{-2T}\mathcal{F}(T)^{2}=\sum_{I\leq 2}\|\nabla^{I}\Sigma\|_{L^{2}(M_{T})}^{2}.$
The third is the pointwise control norm
$\mathcal{N}^{\infty}(T):=\|e^{T}\Sigma\|_{L^{\infty}(M_{T})}+\Bigl\|e^{2T}\Bigl(\frac{N}{n}-1\Bigr)\Bigr\|_{L^{\infty}(M_{T})}.$
The proof is organized around a bootstrap scheme for the triple
$\bigl(\mathcal{O},\mathcal{F},\mathcal{N}^{\infty}\bigr).$
More precisely, on a bootstrap interval $T_{0},T^{\ast})$ we assume
$\mathcal{O}(T)\leq\Gamma,\qquad\mathcal{F}(T)\leq\mathds{Y},\qquad\mathcal{N}^{\infty}(T)\leq\mathds{L},\qquad T\in[T_{0},T^{\ast}),$
with bootstrap constants chosen so that
$(\mathcal{I}^{0})^{2}+\mathcal{I}^{0}+1<\min\{\Gamma,\mathds{Y},\mathds{L}\},\qquad\Gamma+\mathds{Y}+\mathds{L}\leq e^{T_{0}/5},$
where $\mathcal{I}^{0}$ denotes the size of the initial data in the norms relevant to the main theorem. The purpose of the argument is to show that, once $T_{0}=a$ is taken sufficiently large such that $\mathcal{I}^{0}e^{-a/10}<1$, one in fact has the stronger bounds
$\mathcal{O}(T)+\mathcal{F}(T)+\mathcal{N}^{\infty}(T)\lesssim\mathcal{I}^{0}+1\qquad\text{for all }T\in[T_{0},T^{\ast}),$
thereby improving the bootstrap assumptions and extending the solution globally to the future.
The closure mechanism is hierarchical:
$\mathcal{N}^{\infty}\;\Longrightarrow\;\mathcal{F}\;\Longrightarrow\;\mathcal{O}.$
The first implication is obtained by combining the pointwise control of $\Sigma$ and NN with the lower–order evolution equation for $\Sigma$. The decisive term in the corresponding energy identity is the mixed contribution
$\sum_{I\leq 2}e^{-T}\int_{M_{T}}N\,\nabla^{I}\mathfrak{T}\,\nabla^{I}\Sigma\,\mu_{g},$
which satisfies
$\biggl|\sum_{I\leq 2}e^{-T}\int_{M_{T}}N\,\nabla^{I}\mathfrak{T}\,\nabla^{I}\Sigma\,\mu_{g}\biggr|\lesssim e^{-T}\mathcal{O}(T)^{2}.$
Accordingly, one obtains a differential inequality of the form
$\frac{d}{dT}E^{\mathrm{lower}}(T)\leq-2(n-1)E^{\mathrm{lower}}(T)+Ce^{-T}\mathcal{O}(T)^{2}.$
Since $e^{-T}$ is integrable, this forcing term is genuinely lower order in time. After integrating the inequality and using the bootstrap bound for $\mathcal{O}$, one obtains first a uniform estimate for $E^{\mathrm{lower}}$, and then, by a further iteration exploiting the decay already gained, the sharper bound
$\mathcal{F}(T)\lesssim 1+\mathcal{I}^{0}+\mathcal{O}(T)^{2}.$
The key point is that the curvature defect $\mathfrak{T}$ does not need to be small at the initial time; it only needs to be finite. The factor $e^{-T}$ is what converts large but finite geometric input into an integrable error.
The second implication, from $\mathcal{F}$ to $\mathcal{O}$, enters at top order. After commuting the equations and using the elliptic estimates for the lapse, one encounters cubic and quartic expressions in which one differentiated factor is paired with two lower order factors. A representative contribution has the form
$\sum_{I\leq 2}\sum_{m=0}^{I}\sum_{J_{1}+J_{2}+J_{3}=m}e^{-T}\int_{M_{T}}\nabla^{J_{1}}N\,\nabla^{J_{2}+1}\Sigma\,\nabla^{J_{3}+I-m}\Sigma\,\nabla^{I+1}\Sigma\,\mu_{g}.$
Using Sobolev, elliptic, and interpolation estimates, this is bounded by
$\lesssim e^{-2T}\mathcal{F}(T)\mathcal{O}(T)^{2}.$
Once the lower order estimate for $\mathcal{F}$ is available, this becomes
$\lesssim e^{-2T}\mathcal{O}(T)^{4}\lesssim e^{-2T}\Gamma^{4}.$
The coefficient $e^{-2T}$ is now time–integrable with room to spare, and therefore
$\int_{T_{0}}^{T}e^{-2s}\Gamma^{4}\,ds\lesssim e^{-2T_{0}}\Gamma^{4}.$
This is the point at which the “relative smallness” condition in the statement of the main theorem enters: for arbitrarily large but finite initial size $\mathcal{I}^{0}$, one chooses $T_{0}=a$ so large that the integrated top–order errors are perturbative. In other words, the argument does not require absolute smallness of the data; it requires only that the largeness of the data be dominated by the expansion scale present at the initial slice.
Finally, the metric and the gauge must be propagated without loss of geometric control. The metric equation has the schematic form
$\partial_{T}g=e^{-T}N\Sigma-2\Bigl(1-\frac{N}{n}\Bigr)g,$
so once $\mathcal{N}^{\infty}$ is controlled, the deformation of the metric is integrable in time. This yields uniform equivalence of the evolving metrics, control of the volume form and isoperimetric constants, and hence uniform Sobolev inequalities on all future slices. These geometric bounds are then fed back into the elliptic estimates for the lapse and the Sobolev estimates used in the energy argument, closing the system.
[…]
@@
The analytical framework developed herein extends to the case in which the underlying manifold $M$ admits a positive Yamabe invariant, i.e., $\sigma(M)>0$, which includes in particular the setting with positive cosmological constant $\Lambda>0$. Under an additional spectral condition on the Laplace–Beltrami operator associated to the evolving metric, ensuring the uniqueness of the solution to the elliptic lapse equation, we establish convergence of the metric to one with constant positive scalar curvature.
@@
In our context, the underlying cause of non-uniqueness is distinct and originates from the failure of uniqueness of solutions to the lapse equation, rather than the conformal degeneracy of the space of constant scalar curvature metrics. We emphasize that, even in the presence of convergence, the limiting metric is generally not conformal to the initial metric. For completeness, and to illustrate the applicability of our method in this setting, we state the corresponding convergence result below; the proof proceeds with only minor modifications to the arguments presented in the negative Yamabe case. In addition, the construction of the data follows in an exact similar fashion.
@@
(37) $-\Delta_{g}(1-\frac{N}{n})+(|\Sigma|^{2}-\frac{1}{3})(1-\frac{N}{n})=-|\Sigma|^{2},~\Delta_{g}:=g^{ij}\nabla_{i}\nabla_{j}.$
@@
$\Delta_{g}+n\left(|\Sigma|^{2}-\frac{1}{3}\right)$
@@
Theorem 1.2 (Global well-posedness for $\Lambda>0$ and $\sigma(M)>0$).
@@ 1.8. Consequences for Geometrization
An important conceptual motivation underlying the present work concerns the potential role of the Einstein evolution equations in the geometric classification of closed 33-manifolds. Specifically, we consider the relevance of the long-time behavior of solutions to the Einstein flow in connection with Thurston’s geometrization conjecture. In this discussion, we assume throughout that the initial data satisfy the regularity assumptions stipulated in the main theorem.
A central open problem in mathematical general relativity is whether the Einstein equations on globally hyperbolic, spatially compact cosmological spacetimes can dynamically implement a form of geometrization of the underlying spatial manifold. In particular, we consider the case in which the Cauchy hypersurface $M$ is a closed, oriented, connected 33-manifold of negative Yamabe type, i.e., with $\sigma(M)\leq 0$.
@@
Let $M$ be a closed, oriented, connected 33-manifold with $\sigma(M)\leq 0$. A weak geometrization of $M$ is a decomposition
@@
A heuristic explanation for this phenomenon is furnished by comparison with a classical theorem of Cheeger and Gromov [7, 8], which characterizes volume collapse of Riemannian manifolds with bounded curvature. Specifically, if a sequence $\{(M_{j},g_{j})\}$ of compact Riemannian 33-manifolds exhibits collapse with two-sided curvature bounds, then $M_{j}$ admits an $\mathcal{F}$-structure of positive rank, and hence is a graph manifold for sufficiently large $j$. Thus, any dynamical mechanism aiming to realize geometrization via Einstein evolution must, at minimum, produce volume collapse with bounded curvature in the graph component $G\subset M$ for large times.
@@
Let $B(1)\subset M$ denote a geodesic ball of unit radius (scale 11 in injectivity radius scale) with respect to the induced metric at CMC time $T\in[T_{0},\infty)$. Then the volume $\operatorname{Vol}(B(1))_{T}$ remains uniformly bounded above by the initial volume $\operatorname{Vol}(B(1))_{T_{0}}$ for all $T\geq T_{0}$.
@@ 2. Preliminaries
$\Delta_{L}:\Gamma(S^{2}T^{*}M)\to\Gamma(S^{2}T^{*}M),$
@@
Throughout, the spaces of smooth symmetric covariant 2-tensors and vector fields on $M$ are denoted by $S^{0}_{2}(M)$ and $\mathfrak{X}(M)$, respectively. The trace-free and transverse-traceless parts of a symmetric 2-tensor AA are denoted by $\widehat{A}$ and $A^{TT}$, respectively. We will also frequently denote the Cauchy slice at constant mean curvature time $T$ by $M(T)$. In addition, we frequently use the following transport identity, which is a consequence of the transport theorem [32]
@@ 3. Gauge fixed Einstein’s equations
Initial data for 47 consist of an n dimensional manifold $M$, a Riemannian metric $g$, a covariant 2-tensor $k$, all assumed to be smooth and to satisfy
@@
where $k^{TT}$ denotes the transverse-traceless ($T$) part of $k$, satisfying $tr_{g}k^{TT}$ and $\mbox{div }_{g}k^{TT}=0$, and $Z$ is a smooth vector field on $M$.
@@
We show that the non-$T$ part involving the vector field Z must vanish. To that end, consider the $L^{2}$-inner product of the momentum constraint with the vector field Z, and apply Stokes’ theorem:
@@
(62) $R(g)-|k^{\mathrm{TT}}|^{2}=2\Lambda-\tau^{2}\left(1-\frac{1}{3}\right).$
@@
(76) $\mathfrak{T}_{ij}:=\text{Ric}_{ij}-\frac{1}{3}R(g)g_{ij}.$
@@ 3.1. Transported Spatial Coordinates and the Re-Scaled Einstein System
From (79), and the fact that $\widehat{\mathsfbf{n}}=\frac{1}{3}(\partial_{T^{\prime}}-X^{i}\partial_{i})$ in general coordinates, we deduce:
(80) $\mathscr{L}_{\widehat{\mathsfbf{n}}}(x^{i})=\widehat{\mathsfbf{n}}(x^{i})=\frac{1}{3}(\partial_{T^{\prime}}-X^{j}\partial_{j})(x^{i})=-\frac{1}{3}X^{i}\Rightarrow X^{i}=0.$
@@
(83) $-\Delta_{g}\left(\frac{N}{n}-1\right)+\left(|\Sigma|^{2}+\frac{1}{3}\right)\left(\frac{N}{n}-1\right)=-|\Sigma|^{2}.$
@@ 3.2. Integration and Norms
Let $(M,g)$ be a smooth, closed, oriented Riemannian 33-manifold. Fix a smooth partition of unity $\{\xi_{U}\}_{U\in\mathcal{U}}$ subordinate to a finite atlas $\{(U,\varphi_{U})\}_{U\in\mathcal{U}}$ of coordinate charts $\varphi_{U}:U\to\mathbb{R}^{3}$. For a measurable function $f:M\to\mathbb{R}$, the integral of $f$ with respect to the Riemannian volume form $\mu_{g}$ is defined by
@@
We also introduce exponentially weighted norms in time to account for the expanding geometry in cosmological spacetimes. Let $T$ denote the CMC time parameter, and let $1\geq\gamma>0$ be a scaling parameter (which will ultimately be chosen to be $\gamma=1$). For $I\leq 2$, define the weighted Sobolev norms
@@
where NN denotes the spacetime lapse function.